OpenStudy (anonymous):

Determine which equation has a solution of x=1 hath A.5x +8 -12=-6x + 2x - 3 B.3x - 1- 4x=-2x +7x - 4 c.-4x - 6 + x=-2x + x - 10 d.-3x - 8 + x=-4x + x - 10

OpenStudy (anonymous):

just solve each equation and you should get the answer like A 5X+8-12=-6X+2X-3 first combine like terms 5X-4=-4X-3 add 4X to both sides 9X-4=-3 add 4 to both sides 9X=1 divide 9 to both sides 0.11111 and just keep on doing that until you get 1 as an answer

OpenStudy (anonymous):

it's B. 3x - 1 - 4x=-2x + 7x - 4

OpenStudy (anonymous):

yes, correct